Cowling’s rule. Cowling’s rule is another method for
determining the dosage of a drug to prescribe to a child.
For this rule, the formula
d=D(a+1)
24
gives the child’s dosage d, where D is the adult dosage
and a is the age of the child in years. If the adult dosage
of a drug is 600 milligrams and a doctor uses this
formula to determine that a child’s dosage is
200 milligrams, then how old is the child?

Answer by stanbon(75887)

Cowling’s rule. Cowling’s rule is another method for
determining the dosage of a drug to prescribe to a child.
For this rule, the formula
d=D(a+1)
24
---
I am going to assume that 24 is the denominator of the right side.
---
gives the child’s dosage d, where D is the adult dosage
and a is the age of the child in years.
======================================================
If the adult dosage of a drug is 600 milligrams and a doctor uses this
formula to determine that a child’s dosage is 200 milligrams, then how
old is the child?
-----
200 = 600(a+1)/24
Multiplying both sides by 24 you get:
4800 = 600(a+1)
Dividing both sides by 600 you get:
8 = a+1
---
So a = 7 yrs old
======================
Cheers,
Stan H.